OpenStudy (anonymous):

Given f: R->[-1,+1]: x->cos(x) Show f(x) is continuous in its domain with the use of epsilon delta definition

OpenStudy (anonymous):

this is not that easy i think it depends also on your definition of cosine first you prove it is continuous at say 0 and then you prove \[\lim_{h\to 0}\cos(x+h)=\cos(x)\] probably using "addition angle" formula

OpenStudy (anonymous):

you will need a geometric argument, try page 5 and 6 here

OpenStudy (anonymous):

addition angle formula? simpsons formulas?